UT 201206166 AO/PISCES technical night

Observer: DMiller (Tucson), JGuerra, KBoutsia (LBTO)
Telescope Operator: SAllanson (LBTO)
Additional AO Support: RBriguglio, CArcidiacono (LBTO) APuglisi, LFini (Arcetri)
Telescope Support: JUrban(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Tasks completed:

- Switch build to 2012C. Took about an hour because of OSS problems running on the wrong TCS machine
- 7 minutes from preset in ACE mode to closed AO loop. This include "beginning of the night" Act Opt collimation of >4 microns of coma.
- Successful test of offsets while AO loop is closed.
- Collect J, Ks, H best focus. Done

Tasks not completed:

- Have not yet used observe routines on-sky.

Details

During the day, JC installed the webcam to be able to read the Pisces filter wheel encoder. Doug tested communication to Pisces and that new header enriched Pisces images appear in /newdata. There was an issue with the .ssh authorized file because the machine names on the mountain change (lbto.as.arizona.edu -> mountain.lbto.org). Fixed!

During the first half of the night (OSU Science with LBC) JC collected darks at Carmelo's request to be able to determine new corquad parameter

07:00 Sharp! Handover from OSU. Steve reconfigure right side to Gregorian. JC turn off treehouse fans.

07:20 Steve and Dan start change from build 2012F to 2012C

08:20 Build change complete. Had problems with OSS running on TCS1 rather then on TCS3. Dan not sure why this was important.

08:25 Steve set IE/CA and do some collimation by hand. A lot of coma!

08:35 preset to AO674 to close AO loop. Collimated in a few iterations

08:40 runao. Success at 08:42!!

08:42 Send 8 offsets or 0,0, 10,0, 0,0, 2,0, 10,0, 0,0 400,0 0,0, 0,10 0,400 10,400 All successful! For those outside the patrol range of Wunit and AGw, the probes did not move. When we offset back inside the patrol range, the both moved to the correct position.

08:59 GCS complain that images not arriving. Steve will reset azcam server on the next preset

09:00 Alfio and JC are checking AO performance. Getting ready to determine best focus on Pisces.

09:05 images pisces 5-7 look like the start is a triple star. Reflections?

09:10 right_wfscimage 50-60 will good to make a wfsc reference vector for Active Optics. AO loop is closed.

09:25 Alfio have the same .ssh/authorized_keys file for wfsdx to LBTO@obs3. Fixed. Ready to find best focus with AutoFocusPisces.py

09:50 Start auto focus routines. A few problems

10:20 Finished with first set to focus data. The starting focus value was too far from best focus, so all data were on once side of the caustic. The starting focus position was changed and another set of focus data started

10:30 Start second set of J focus data

10:45 Determine J best focus to be 51.4 mm from 51.1 mm. JC put this in a config file

10:47 Start collecting Ks best focus data. 52.3 mm, same as before

11:05 Start collecting H best focus data. 51.7 mm, same as before

11:35 Done with focus. Move to Zenith, collect flats.

-- DougMiller - 16 Jun 2012
Topic revision: r5 - 16 Jun 2012, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback